Mystic Beverage Company is considering purchasing a new bottling machine. The new machine costs $130,000, plus...

  1. Mystic Beverage Company is considering purchasing a new bottling machine. The new machine costs $130,000, plus installation fees of $10,000 and will generate earning before interest and taxes of $60,000 per year over its 6-year life. The machine will be depreciated on a straight-line basis over its 6-year life to an estimated salvage value of 0. Mystic’s marginal tax rate is 40%. Mystic will require $50,000 in NWC if the machine is purchased.

(a) determine the initial outlay

(b) determine the annual cash flows in years 1-6

(c) determine the terminal cash flow
Answer #1

a: Initial outlay = cost + installation + working capital = 130000+10000+50000

= 190000

b: Annual cash flows = EBIT*(1-tax)+Tax*depreciation

= 60000*(1-40%)+40%*140000/6

= 45333.33

c: Terminal cash flow = Working capital recovery + OCF in year 6

=50000+45333.33

= 95333.33